# On Probabilistic Pullback Metrics on Latent Hyperbolic Manifolds

### v1

@@ Abstract
Gaussian Process Latent Variable Models (GPLVMs) have proven effective in capturing complex, high-dimensional data through lower-dimensional representations. Recent advances show that using Riemannian manifolds as latent spaces provides more flexibility to learn higher quality embeddings. This paper focuses on the hyperbolic manifold, a particularly suitable choice for modeling hierarchical relationships. While previous approaches relied on hyperbolic geodesics for interpolating the latent space, this often results in paths crossing low-data regions, leading to highly uncertain predictions. Instead, we propose augmenting the hyperbolic metric with a pullback metric to account for distortions introduced by the GPLVM’s nonlinear mapping. Through various experiments, we demonstrate that geodesics on the pullback metric not only respect the geometry of the hyperbolic latent space but also align with the underlying data distribution, significantly reducing uncertainty in predictions.
@@ 1 Introduction
Hyperbolic geometry is particularly useful in fields where data exhibits clear hierarchical structures, such as natural language processing for representing word hierarchies and taxonomies (Nickel and Kiela,, 2017, 2018), as well as in social network analysis for modeling community structures (Krioukov et al.,, 2010; Doorenbos et al.,, 2024). Additionally, it finds applications in biology (Alanis-Lobato et al.,, 2018), human motion taxonomies (Jaquier et al.,, 2024), and computer vision (Khrulkov et al.,, 2020). However, many datasets in these disciplines are high dimensional and hetereogenous, making latent variable models (LVMs) indispensable. Recently, hyperbolic embeddings gained significant attention due to their ability to capture hierarchical structures inherent in complex high-dimensional data (Skopek et al.,, 2020; Cho et al.,, 2023; Jaquier et al.,, 2024). These embeddings leverage hyperbolic growth to accommodate hierarchical relationships, present in trees or graphs, that are difficult to model in Euclidean spaces (Cvetkovski and Crovella,, 2009).
@@
Despite the potential of hyperbolic embeddings, most of the state-of-the-art techniques operate without accounting for the intrinsic properties of the data. For example, the learned embeddings should be distance-preserving and their distribution should closely match that of the observed data. By doing so, any operation on the hyperbolic latent space complies with the properties of the data manifold. This problem can be understood from a differential-geometry point-of-view (Hauberg,, 2019). For instance, recent works have focused on integrating the data manifold structure into latent spaces through stochastic pullback metrics (Tosi et al.,, 2014; Arvanitidis et al.,, 2018, 2021). This approach not only introduces uncertainty information into the latent space via the Riemannian metric, but also enables the generation of geodesics that more closely align with the true data distribution. The latter has been leveraged in robot motion generation (Beik-Mohammadi et al.,, 2021), protein sequencing (Detlefsen et al.,, 2022), and data augmentation for medical imaging (Chadebec et al.,, 2023). However, none of the aforementioned works considered hyperbolic geometry as inductive bias on the latent space.
@@
This paper makes several key contributions to the field of hyperbolic latent variable models and GPLVMs. First, we introduce a general formulation of the Riemannian pullback metric on hyperbolic latent spaces, under the assumption of a stochastic latent-to-ambient mapping. Second, we present a novel development of the Riemannian pullback metric tailored to Gaussian Process Hyperbolic Latent Variable Models (GPHLVMs) (Jaquier et al.,, 2024), incorporating appropriate Riemannian projections onto tangent spaces to account for the hyperbolic geometry. Third, we develop the kernel derivatives within this setting, highlighting the limitations of current autodifferentiation techniques when applied to our setting. Finally, we demonstrate the benefits of hyperbolic pullback-based geodesics on four experiments: A proof-of-concept $\mathsf{C}$-shape example, MNIST data interpolation, multi-cellular robot design (Dong et al.,, 2024), and human grasps generation (Jaquier et al.,, 2024).
@@ 2 Background
Before delving into the hyperbolic manifold, it is necessary to establish a basic understanding in Riemannian geometry (Lee,, 2018). A Riemannian manifold $\mathcal{M}$ is a smooth manifold equipped with a smoothly varying inner product ${g_{\bm{x}}\!:\mathcal{T}_{\bm{x}}\mathcal{M}\!\times\!\mathcal{T}_{\bm{x}}\mathcal{M}\!\rightarrow\!\mathbb{R}:(\bm{u},\bm{v})\!\mapsto\!\langle\bm{u},\bm{v}\rangle_{\bm{x}}\!=\!\bm{u}^{\mathsf{T}}\bm{G}_{\bm{x}}\bm{v}}$. The inner product $g_{\bm{x}}$ and equivalently the matrix $\bm{G}_{\bm{x}}$ are known as the Riemannian metric or metric tensor at each point $\bm{x}\!\in\!\mathcal{M}$. The tangent space $\mathcal{T}_{\bm{x}}\mathcal{M}$ is the Euclidean space of tangent vectors of all the possible smooth curves passing through $\bm{x}$. To operate with Riemannian manifolds, it is common practice to exploit the Euclidean tangent spaces and shortest-path curves, so-called geodesics, between two points $\bm{x},\bm{y}\in\mathcal{M}$. The exponential map $\text{Exp}_{\bm{x}}(\bm{u})=\bm{y}$ maps a tangent space vector $\bm{u}\in\mathcal{T}_{\bm{x}}\mathcal{M}$ to a point $\bm{y}$ on the manifold, so that $\bm{y}$ lies on a geodesic starting at $\bm{x}$ in the direction $\bm{u}$, and such that the geodesic distance $\text{d}_{\mathcal{M}}(\bm{x},\bm{y})$ between $\bm{x}$ and $\bm{y}$ equals the length of $\bm{u}$ given by the Riemannian norm $\|\bm{u}\|_{\bm{x}}=\sqrt{\langle\bm{u},\bm{u}\rangle_{\bm{x}}}$. The inverse operation is the logarithmic map $\text{Log}_{\bm{x}}(\bm{y})=\bm{u}$. Finally, the parallel transport $\Gamma_{\bm{x}\rightarrow\bm{y}}(\bm{u})=\bm{v}$ moves a tangent space vector $\bm{u}$ to the tangent space $\mathcal{T}_{\bm{y}}\mathcal{M}$ while preserving the Riemannian inner product.
The hyperbolic manifold is the only Riemannian manifold with constant negative curvature (Ratcliffe,, 2019). It is commonly represented by either the Poincaré model $\mathbb{H}^{D}_{\mathcal{P}}$ (Poincaré,, 1900), employing local coordinates within the unit sphere, or the Lorentz model $\mathbb{H}^{D}_{\mathcal{L}}$ (Jansen,, 1909; Reynolds,, 1993), which uses Cartesian coordinates to represent the surface embedded in $\mathbb{R}^{D+1}$. In this paper we mostly rely on the latter,
since it is numerically more stable. Further details on the hyperbolic manifold are provided in App. A.
@@
A GPLVM generates observations ${\bm{Y}=[\bm{y}_{1}\ldots\bm{y}_{N}]^{\mathsf{T}}\in\mathbb{R}^{N\times D_{y}}}$ from latent variables ${\bm{X}\!=\![\bm{x}_{1}\ldots\bm{x}_{N}]^{\mathsf{T}}\!\in\!\mathbb{R}^{N\times D_{x}}}$ with $D_{x}<D_{y}$ through a non-linear transformation modeled by Gaussian processes (GPs) (Lawrence,, 2003). A GPLVM is formally described by,
$p(\bm{Y}\mid\bm{X},\bm{\Theta})=\prod_{d=1}^{D_{y}}\mathcal{N}(\bm{Y}_{d}\mid\bm{0},\bm{K}_{X}+\sigma_{y}^{2}\bm{I}_{N})\,,$ (2)
where $\bm{Y}_{d}\in\mathbb{R}^{N}$ is the $d$-th observation column, ${\bm{K}_{X}\in\mathbb{R}^{N\times N}}$ is the kernel matrix $(\bm{K}_{X})_{ij}=k(\bm{x}_{i},\bm{x}_{j})$, and $\bm{\Theta}=\{\tau,\kappa,\sigma_{y}^{2}\}$ is the set of scalar hyperparameters, namely, the kernel variance $\tau$, the lengthscale $\kappa$, and the noise variance $\sigma_{y}^{2}$.
When the latent space is Euclidean, a common choice of kernel function is the Squared Exponential (SE) kernel $k^{\text{SE}}(\bm{x},\bm{x}^{\prime})=\tau\exp\left(-\frac{\|\bm{x}-\bm{x}^{\prime}\|^{2}}{2\kappa^{2}}\right)$. As discussed by Jaquier et al., (2024), an Euclidean latent space may not necessarily comply with structural biases in the observed data, and therefore curved geometries such as the hyperbolic manifold may be preferred. In such case the latent variables $\bm{x}_{n}\in\mathbb{H}_{\mathcal{L}}^{D_{x}}$ live in hyperbolic space, and consequently the kernel function is replaced by a hyperbolic kernel as introduced in the next section. Moreover, the latent variables are assigned a hyperbolic wrapped Gaussian distribution as proposed by (Jaquier et al.,, 2024), thus inducing the prior $p(\bm{X}\mid\alpha)=\prod_{n=1}^{N}\mathcal{N}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x}_{n}\mid\bm{\mu}_{0},\alpha\bm{I}_{D_{x}})$, with $\alpha$ controlling the spread of the latent variables (see App. A for more details). The GPHLVM latent points and hyperparameters are inferred via,
$\operatorname*{argmax}_{\bm{X},\bm{\Theta},\alpha}\,p(\bm{Y}\mid\bm{X},\Theta)\,p(\bm{X}\mid\alpha),$ (3)
which can be solved by finding a MAP estimate or via variational inference similar as in the Euclidean case.
@@
The SE and Matérn kernels are standard choices when designing Gaussian Processes (Rasmussen and Williams,, 2006). These kernels have been recently generalized to non-Euclidean spaces such as manifolds (Borovitskiy et al.,, 2020; Jaquier et al.,, 2021), or graphs (Borovitskiy et al.,, 2021).Grigoryan and Noguchi, (1998) demonstrated that, in hyperbolic space, two- and three-dimensional heat kernels suffice since higher-dimensional kernels can be expressed as derivatives of these lower-dimensional ones. The two- and three-dimensional hyperbolic heat kernels are given as,
@@
$k^{\mathbb{H}^{2}_{\mathcal{L}}}(\bm{x},\bm{x}^{\prime})$ $=\dfrac{\tau}{C_{\infty}}\int_{\rho}^{\infty}\dfrac{s\,e^{-\frac{s^{2}}{2\kappa^{2}}}}{\sqrt{\text{cosh}(s)-\text{cosh}(\rho)}}\text{d}s\,,$ (4)
@@
$k^{\mathbb{H}^{3}_{\mathcal{L}}}(\bm{x},\bm{x}^{\prime})$ $=\dfrac{\tau}{C_{\infty}}\dfrac{\rho}{\text{sinh}(\rho)}e^{-\frac{\rho^{2}}{2\kappa^{2}}}\,,$ (5)
@@
with points on the Lorentz model $\bm{x},\bm{x}^{\prime}\in\mathbb{H}^{D_{x}}_{\mathcal{L}}$, their geodesic distance $\rho=\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x},\bm{x}^{\prime})$, kernel variance and lengthscale parameters $\tau,\kappa\in\mathbb{R}_{+}$, and a normalization constant $C_{\infty}=(4\pi\kappa^{2})^{\frac{3}{2}}$. In the 2D case no closed form solution is known. Instead, the kernel needs to be approximated via a discretization of the integral in (4). We use the Monte Carlo approximation introduced by Jaquier et al., (2024),
$k^{\mathbb{H}^{2}_{\mathcal{L}}}(\bm{x},\bm{x}^{\prime})\!\approx\dfrac{\tau}{C_{\infty}}\dfrac{1}{L}\sum_{l=1}^{L}s_{l}\text{tanh}(\pi s_{l})\Phi_{l}(\bm{x}_{\mathcal{P}},\bm{x}^{\prime}_{\mathcal{P}})\,,$ (6)
@@
where $\Phi_{l}(\bm{x}_{\mathcal{P}},\bm{x}^{\prime}_{\mathcal{P}})=\phi_{l}(\bm{x}_{\mathcal{P}})\bar{\phi}_{l}(\bm{x}^{\prime}_{\mathcal{P}})$, $\bar{\phi}_{l}$ denotes the complex conjugate of $\phi_{l}(\bm{x}_{\mathcal{P}})=e^{(1+2s_{l}i)\langle\bm{x}_{\mathcal{P}},\bm{b}_{l}\rangle}$, ${s_{l}\sim e^{-\frac{s^{2}\kappa^{2}}{2}}1_{[0,\infty]}(s)}$ are samples from a truncated Gaussian distribution, $\bm{b}_{l}\sim U(\mathbb{S}^{2})$ are samples from the unit circle, and $\langle\bm{x}_{\mathcal{P}},\bm{b}\rangle=\frac{1}{2}\text{log}\left(\frac{1-|\bm{x}_{\mathcal{P}}|^{2}}{|\bm{x}_{\mathcal{P}}-\bm{b}|^{2}}\right)$ is the hyperbolic outer product with $\bm{x}_{\mathcal{P}}\!=\!\frac{1}{1+x_{0}}[x_{1}\ldots x_{D}]^{\mathsf{T}}\in\mathbb{H}^{D_{x}}_{\mathcal{P}}$ being the Poincaré representation of $\bm{x}$.

### v2

@@ Abstract
Probabilistic Latent Variable Models (LVMs) have proven effective in capturing complex, high-dimensional data through lower-dimensional representations. Recent advances show that using Riemannian manifolds as latent spaces provides more flexibility to learn higher quality embeddings. This paper focuses on the hyperbolic manifold, a particularly suitable choice for modeling hierarchical relationships. Previous approaches relying on hyperbolic geodesics for interpolating the latent space often generate paths crossing low-data regions, leading to highly uncertain predictions. Instead, we propose augmenting the hyperbolic metric with a pullback metric to account for distortions introduced by the LVM’s nonlinear mapping and provide a complete development for pullback metrics of Gaussian Process LVMs (GPLVMs). Our experiments demonstrate that geodesics on the pullback metric not only respect the geometry of the hyperbolic latent space but also align with the underlying data distribution, significantly reducing uncertainty in predictions.
Machine Learning, ICML
[…]
figurec
@@ 1 Introduction
Hyperbolic geometry is particularly useful in fields where data exhibits clear hierarchical structures, such as natural language processing for representing word hierarchies and taxonomies (Nickel & Kiela, 2017, 2018), as well as in social network analysis for modeling community structures (Krioukov et al., 2010; Doorenbos et al., 2024). Additionally, it finds applications in biology (Alanis-Lobato et al., 2018), human motion taxonomies (Jaquier et al., 2024), and computer vision (Khrulkov et al., 2020). However, many datasets in these disciplines are high dimensional and hetereogenous, making latent variable models (LVMs) indispensable. Recently, hyperbolic embeddings gained significant attention due to their ability to capture hierarchical structures inherent in complex high-dimensional data (Skopek et al., 2020; Cho et al., 2023; Yang et al., 2023). These embeddings leverage hyperbolic growth to accommodate hierarchical relationships, present in trees or graphs, that are difficult to model in Euclidean spaces (Cvetkovski & Crovella, 2009). Hyperbolic embeddings excel at preserving the hierarchical relationships of the original data in low-dimensional spaces (Sala et al., 2018). This can be achieved by leveraging a known taxonomy to guide the embedding process (Nickel & Kiela, 2017; Jaquier et al., 2024), by estimating the original data hierarchy via diffusion geometry (Lin et al., 2023), or by learning a tree on the original data (Sonthalia & Gilbert, 2020).
Figure 1: Hyperbolic pullback metric on the tangent space of the Lorentz model. The corresponding pullback geodesic ( ) follow the data ( ) manifold, in contrast to the hyperbolic geodesic ( ).
@@
Despite the potential of hyperbolic embeddings, most of the state-of-the-art techniques operate without accounting for the intrinsic properties of the data. For example, the learned embeddings should be distance-preserving and their distribution should closely match that of the observed data. By doing so, any operation on the hyperbolic latent space complies with the properties of the data manifold. Instead, existing techniques generate geodesics that often traverse low-density regions of the data space, thus resulting in unrealistic interpolations and inaccurate representations. This problem of integrating the data manifold structure into latent spaces can be understood from a differential-geometry point-of-view (Hauberg, 2019) through stochastic pullback metrics (Tosi et al., 2014; Arvanitidis et al., 2018, 2021). This approach not only introduces uncertainty information into the latent space via the Riemannian metric, but also enables the generation of geodesics that faithfully follow the true data distribution. It has been leveraged in robot motion generation (Beik-Mohammadi et al., 2021), protein sequencing (Detlefsen et al., 2022), and data augmentation for medical imaging (Chadebec et al., 2023). However, none of the aforementioned works considered hyperbolic geometry as inductive bias on the latent space.
@@
This paper makes several key contributions to the field of hyperbolic LVMs and GPLVMs. First, we introduce a general formulation of the Riemannian pullback metric on hyperbolic latent spaces, under the assumption of a stochastic latent-to-ambient mapping (see Fig. 1). Second, we present a novel development of the Riemannian pullback metric tailored to Gaussian Process Hyperbolic Latent Variable Models (GPHLVMs) (Jaquier et al., 2024), incorporating appropriate Riemannian projections onto tangent spaces to account for the hyperbolic geometry. Third, we develop the kernel derivatives within this setting, highlighting the limitations of current autodifferentiation techniques when applied to our setting. Finally, we demonstrate the benefits of hyperbolic pullback-based geodesics on four experiments: A proof-of-concept $\mathsf{C}$-shape example, MNIST data interpolation, multi-cellular robot design (Dong et al., 2024), and human grasps generation (Jaquier et al., 2024).
@@ 2 Background
Riemannian geometry:
Before delving into the hyperbolic manifold, it is necessary to establish a basic understanding in Riemannian geometry (Lee, 2018). A Riemannian manifold $\mathcal{M}$ is a smooth manifold equipped with a Riemannian metric, i.e., a smoothly-varying inner product ${g_{\bm{x}}\!:\mathcal{T}_{\bm{x}}\mathcal{M}\!\times\!\mathcal{T}_{\bm{x}}\mathcal{M}\!\rightarrow\!\mathbb{R}}$ over tangent spaces $\mathcal{T}_{\bm{x}}\mathcal{M}$. When considering coordinates, the Riemannian metric is represented in matrix form for each $\bm{x}\!\in\!\mathcal{M}$ as $\langle\bm{u},\bm{v}\rangle_{\bm{x}}\!=\!\bm{u}^{\mathsf{T}}\bm{G}_{\bm{x}}\bm{v}$ with $\bm{u},\bm{v}\in\mathcal{T}_{\bm{x}}\mathcal{M}$. The Riemannian metric defines the length of curves in $\mathcal{M}$, leading to the notion of geodesics, defined as locally length-minimizing curves. To operate with Riemannian manifolds, it is common practice to exploit the Euclidean tangent spaces and the geodesics. The exponential map $\text{Exp}_{\bm{x}}(\bm{u})\!=\!\bm{y}$ maps $\bm{u}\in\mathcal{T}_{\bm{x}}\mathcal{M}$ to a point $\bm{y}\in\mathcal{M}$, so that $\bm{y}$ lies on a geodesic starting at $\bm{x}$ in the direction $\bm{u}$, and such that the geodesic distance $\text{d}_{\mathcal{M}}(\bm{x},\bm{y})$ equals the length of $\bm{u}$ given by $\|\bm{u}\|_{\bm{x}}\!=\!\sqrt{\langle\bm{u},\bm{u}\rangle_{\bm{x}}}$. The inverse operation is the logarithmic map $\text{Log}_{\bm{x}}(\bm{y})\!=\!\bm{u}$. Finally, the parallel transport $\Gamma_{\bm{x}\rightarrow\bm{y}}(\bm{u})=\bm{v}$ moves a vector $\bm{u}\in\mathcal{T}_{\bm{x}}\mathcal{M}$ to $\mathcal{T}_{\bm{y}}\mathcal{M}$ while preserving the Riemannian inner product.
Optimizing manifold-valued functions requires generalizing the notion of gradient. The Riemannian gradient $\mathrm{grad}_{\bm{x}}(f)$ of a function $f:\mathcal{M}\to\mathbb{R}$ at $\bm{x}\in\mathcal{M}$ is the unique tangent vector in $\mathcal{T}_{\bm{x}}\mathcal{M}$ that satisfies $\mathcal{D}_{\bm{u}}f(\bm{x})=\langle\mathrm{grad}_{\bm{x}}f(\bm{x}),\bm{u}\rangle_{\bm{x}}$, where $\mathcal{D}_{\bm{u}}f(\bm{x})$ is the directional derivative of $f$ in the direction $\bm{u}\in\mathcal{T}_{\bm{x}}\mathcal{M}$ (Boumal, 2023, Chap. 3). The Riemannian Jacobian $\tilde{\bm{J}}$ of a function $f:\mathcal{M}\to\mathbb{R}^{D}$ is composed by the Riemannian gradients for each output dimension, i.e.,
$\tilde{\bm{J}}=[\mathrm{grad}_{\bm{x}}(f_{1})\>\ldots\>\mathrm{grad}_{\bm{x}}(f_{D})]^{\mathsf{T}}.$ (1)
@@
The hyperbolic manifold is the only Riemannian manifold with constant negative curvature (Ratcliffe, 2019). It is commonly represented by either the Poincaré model $\mathbb{H}^{D}_{\mathcal{P}}$ (Poincaré, 1900), employing local coordinates within the unit ball, or the Lorentz model $\mathbb{H}^{D}_{\mathcal{L}}$ (Jansen, 1909; Reynolds, 1993), using Cartesian coordinates to represent the surface embedded in $\mathbb{R}^{D+1}$. In this paper we mostly rely on the latter, which is numerically more stable and defined as,
$\mathbb{H}^{D}_{\mathcal{L}}=\{\bm{x}\in\mathbb{R}^{D+1}\mid\langle\bm{x},\bm{x}\rangle_{\mathcal{L}}=-1,x_{0}>0\},$ (2)
where $\langle\bm{x},\bm{y}\rangle_{\mathcal{L}}=\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}\bm{y}$ is the Lorentzian inner product with $\bm{G}^{\mathcal{L}}=\text{diag}(-1,1,...,1)$. Further details on the hyperbolic manifold and its operations are provided in App. A.
@@
A GPLVM defines a generative mapping from latent variables $\bm{x}_{n}\!\in\!\mathbb{R}^{D_{x}}$ to observations $\bm{y}_{n}\!\in\!\mathbb{R}^{D_{y}}$ with $D_{x}<D_{y}$ through a non-linear transformation modeled by Gaussian processes (GPs) (Lawrence, 2003). A GPLVM assumes that the observations are normally distributed, i.e., $y_{n,d}\sim\mathcal{N}(f_{n,d},\sigma_{y}^{2})$, and described via a GP and a prior,
$f_{n,d}\sim\text{GP}(0,k(\bm{x}_{n},\bm{x}_{n}))\quad\text{and}\quad\bm{x}_{n}\sim\mathcal{N}(\bm{0},\bm{I}),$ (3)
where $y_{n,d}$ is the $d$-th dimension of $\bm{y}_{n}$, $k:\mathbb{R}^{D_{x}}\times\mathbb{R}^{D_{x}}\to\mathbb{R}$ is the GP kernel function which measures the similarity between two inputs $\bm{x}_{n},\bm{x}_{m}$, and $\sigma_{y}^{2}$ is the noise variance.
In some cases, a Euclidean latent space may not necessarily comply with structural biases in the observed data, and therefore curved geometries such as the hyperbolic manifold may be preferred. In such cases, the latent variables $\bm{x}_{n}\in\mathbb{H}_{\mathcal{L}}^{D_{x}}$ live in a hyperbolic space and the generative mapping is defined via a GPHLVM (Jaquier et al., 2024). Consequently the kernel function in (3) is replaced by a hyperbolic kernel $k^{\mathbb{H}_{\mathcal{L}}^{D_{x}}}:\mathbb{H}_{\mathcal{L}}^{D_{x}}\times\mathbb{H}_{\mathcal{L}}^{D_{x}}\to\mathbb{R}$ as introduced shortly. Moreover, the latent variables are assigned a hyperbolic wrapped Gaussian prior $\bm{x}_{n}\sim\mathcal{N}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{\mu}_{0},\alpha\bm{I})$, with $\bm{\mu}_{0}=(1,0,\ldots,0)^{\mathsf{T}}$ and $\alpha$ controlling the spread of the latent variables (see App. A for more details). The GPHLVM latent variables and hyperparameters are inferred by finding a MAP estimate or via variational inference similar as in the Euclidean case.
@@
The SE and Matérn kernels are standard choices when designing GPs (Rasmussen & Williams, 2006). These kernels have been recently generalized to non-Euclidean spaces such as manifolds (Borovitskiy et al., 2020; Jaquier et al., 2021), or graphs (Borovitskiy et al., 2021). Grigoryan & Noguchi (1998) demonstrated that, in hyperbolic space, $2$- and $3$-dimensional SE kernels suffice since higher-dimensional kernels can be expressed as derivatives of these lower-dimensional ones. The 2D and 3D hyperbolic SE kernels are given as,
@@
$k^{\mathbb{H}^{2}_{\mathcal{L}}}(\bm{x},\bm{z})$ $=\dfrac{\tau}{C_{\infty}}\int_{\rho}^{\infty}\dfrac{s\,e^{-\frac{s^{2}}{2\kappa^{2}}}}{\sqrt{\text{cosh}(s)-\text{cosh}(\rho)}}\text{d}s\,,$ (4)
@@
$k^{\mathbb{H}^{3}_{\mathcal{L}}}(\bm{x},\bm{z})$ $=\dfrac{\tau}{C_{\infty}}\dfrac{\rho}{\text{sinh}(\rho)}e^{-\frac{\rho^{2}}{2\kappa^{2}}}\,,$ (5)
@@
where $\rho=\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x},\bm{z})$ is the geodesic distance between $\bm{x},\bm{z}\in\mathbb{H}^{D_{x}}_{\mathcal{L}}$, $\tau,\kappa\in\mathbb{R}_{+}$ are the kernel variance and lengthscale, and $C_{\infty}$ is a normalization constant. As no closed form solution is known in the $2$D case, the kernel needs to be approximated via a discretization of the integral in (4). We use the positive semidefinite Monte Carlo approximation introduced by Jaquier et al. (2024),
$k^{\mathbb{H}^{2}_{\mathcal{L}}}(\bm{x},\bm{z})\!\approx\dfrac{\tau}{C_{\infty}}\dfrac{1}{L}\sum_{l=1}^{L}s_{l}\text{tanh}(\pi s_{l})\Phi_{l}(\bm{x}_{\mathcal{P}},\bm{z}_{\mathcal{P}})\,,$ (6)
@@
where $\Phi_{l}(\bm{x}_{\mathcal{P}},\bm{z}_{\mathcal{P}})\!=\!\phi_{l}(\bm{x}_{\mathcal{P}})\bar{\phi}_{l}(\bm{z}_{\mathcal{P}})$ is the only component depending on the kernel inputs, $\bar{\phi}_{l}$ is the complex conjugate of $\phi_{l}(\bm{x}_{\mathcal{P}})\!=\!e^{(1+2s_{l}i)\langle\bm{x}_{\mathcal{P}},\bm{b}_{l}\rangle}$, $\bm{x}_{\mathcal{P}}\in\mathbb{H}^{D_{x}}_{\mathcal{P}}$ is the Poincaré representation of $\bm{x}$, $\bm{b}_{l}\sim U(\mathbb{S}^{1})$ are samples from the unit circle, ${s_{l}\sim e^{-\frac{s^{2}\kappa^{2}}{2}}1_{[0,\infty]}(s)}$ are samples from a truncated Gaussian distribution, and $\langle\bm{x}_{\mathcal{P}},\bm{b}\rangle\!=\!\frac{1}{2}\text{log}\left(\frac{1-|\bm{x}_{\mathcal{P}}|^{2}}{|\bm{x}_{\mathcal{P}}-\bm{b}|^{2}}\right)$ is the hyperbolic outer product.
Pullback metrics:
Consider an immersion ${f\colon\mathcal{S}\to\mathcal{M}}$ from a latent space $\mathcal{S}$ to a Riemannian manifold $\mathcal{M}$ equipped with a Riemannian metric $g_{\bm{y}}$. The immersion $f$ induces a pullback metric $g_{\bm{x}}^{\text{P}}$ on $\mathcal{S}$ which, for $\bm{x}\in\mathcal{S}$ and $\bm{u},\bm{v}\in\mathcal{T}_{\bm{x}}\mathcal{S}$, is given by (Lee, 2018, Chap. 2)
$g^{\text{P}}_{\bm{x}}(\bm{u},\bm{v})=g_{f(\bm{x})}\big{(}df_{\bm{x}}(\bm{u}),df_{\bm{x}}(\bm{v})\big{)}.$ (7)
With coordinates, the pullback metric is given in matrix form by $\bm{G}^{\text{P}}_{\bm{x}}=\tilde{\bm{J}}^{\mathsf{T}}\bm{G}_{\bm{y}}\tilde{\bm{J}}$, where $\tilde{\bm{J}}$ is the Riemannian Jacobian (1) of $f$ at $\bm{x}$. Intuitively, $g^{\text{P}}_{\bm{x}}$ evaluates on tangent vectors of $\mathcal{T}_{\bm{x}}\mathcal{S}$ by moving them to $\mathcal{T}_{f(\bm{x})}\mathcal{M}$ to compute their inner product. For an immersion $f:\mathcal{S}\to\mathbb{R}^{D_{y}}$ with Euclidean co-domain, i.e., $\bm{G}_{\bm{y}}=\bm{I}$, the pullback metric is $\bm{G}^{\text{P}}_{\bm{x}}=\tilde{\bm{J}}^{\mathsf{T}}\tilde{\bm{J}}$. For Euclidean domains $\mathcal{S}=\mathbb{R}^{D_{x}}$, $\tilde{\bm{J}}$ equals the Euclidean Jacobian $\bm{J}=[\frac{\partial f_{1}}{\partial\bm{x}}\>\ldots\>\frac{\partial f_{D_{y}}}{\partial\bm{x}}]^{\mathsf{T}}\in\mathbb{R}^{D_{y}\times D_{x}}$. In this paper, we define pullback metrics of hyperbolic LVMs by explicitly leveraging immersions $f\colon\mathbb{H}^{D_{x}}\to\mathbb{R}^{D_{y}}$, where $f$ is a GPHLVM.